OpenStudy (anonymous):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (anonymous):

@bananaroof01

OpenStudy (anonymous):

@doddododo

jimthompson5910 (jim_thompson5910):

@helper9000 is it possible to pick 2 numbers from {0, 2, 5, 8}, add them up, and get something that isn't in the set? if so, which two values did you pick?

OpenStudy (anonymous):

hi

OpenStudy (mrnood):

@helper9000 when you have answered @jim_thompson5910 question about the first set , then ask it about the other three options. A set is 'closed' under addition if, when you add ANY 2 members of the set , the ne the result IS always a member of the set

OpenStudy (anonymous):

hi and there is only one that dose not add to an even number and it is 8+5=13

OpenStudy (anonymous):

i think its b

OpenStudy (mrnood):

@doddododo that answer is not helpful as it does not explain - it is also wrong

OpenStudy (anonymous):

story

OpenStudy (mrnood):

@helper9000 your last reply does not answer the question that @jim_thompson5910 asked

OpenStudy (anonymous):

The only ones is I see that is closed under addition is D, C and B because A. 5+8 = 13 and 13 is not in the set so A is out B. say 2 and 4 for example 2+4 = 6 all even integers C. Same for rational number example 4 + 3 + 7 which is a rational number D 0+0=0 so that is closed under addition too

OpenStudy (ninjaslice):

So, hang on. you want us to do you work for you? Isn't that like uh, doing a test for you bruh? @helper9000

OpenStudy (anonymous):

oh sorry i read it wrong

jimthompson5910 (jim_thompson5910):

@helper9000 you are correct in saying that 5+8 = 13 which is NOT in the set but we don't need to worry about if the result is even or not so because 13 isn't in the set, this means the set in choice A is NOT closed under addition

OpenStudy (anonymous):

no i just want other people to help me out and make sure it is right

OpenStudy (ninjaslice):

People, who even cared to read the privacy policy? NO KIDS UNDER 13

OpenStudy (anonymous):

ok so why dose that matter?

OpenStudy (ninjaslice):

@helper9000 Well you see, when people Like I don't know who but i seem to b writing a topic on his/her topic..... It seems to me that it is fairly important, due to you have people who want the answer, but do not want it to be explained. such as other kids who ask what is 5+5 =? so come on man, google is on your side here, not this site.....

OpenStudy (anonymous):

whats 5+5=

OpenStudy (ninjaslice):

@doddododo .......... Your joking right? yes you are.

OpenStudy (anonymous):

yes i was just kidding don't un fan my pleas

OpenStudy (mrnood):

@ninjaslice the rules here say that helpers should not give complete answers , but shpuld guide the poster to work out the answer for themselves (just like your teacher does not say 'google it') however @Alina_Dubrovnik answer seems to give allthe answers and not leave any work for the poster to do. @doddododo your answer was wrong only in as much as it is incomplete

OpenStudy (anonymous):

story I meant don't UN fan me

OpenStudy (ninjaslice):

@MrNood your wrong, because my science teacher actually told me to google my own SCIENCE HOMEWORK, know facts before you try to sound better.

OpenStudy (mrnood):

my point is that this site IS for help with people's work - but should be an aid to learning, not simply a source of direct answers

OpenStudy (ninjaslice):

@MrNood That is my point as well. I do not want to just be like OH OKAY SO.... 5+5 EQUALS 10 LETS MOVE ON NOW.

OpenStudy (ninjaslice):

Good we have our own facts right.

OpenStudy (mrnood):

and btw - that should be "you're wrong" not "your wrong"

OpenStudy (anonymous):

MrNood fan me please

OpenStudy (anonymous):

i fned you

OpenStudy (mrnood):

fan you for what?

OpenStudy (anonymous):

i faned you

OpenStudy (ninjaslice):

@MrNood Yes, I am highly aware of that mistake. Due to I wanted to see if you were a Grammar Nancy. I did that on purpose. You are acting like you know it all. So if you know it all feel free to 100% explain the question above and lets stop this argument. P.S I am not dumb.

OpenStudy (mrnood):

PS you did it twice - on an earlier post too

OpenStudy (ninjaslice):

@MrNood Yes I am aware :)

OpenStudy (ninjaslice):

@MrNood Btw, I do not see any periods after your sentence so.....

OpenStudy (mrnood):

i'm out of here!

OpenStudy (ninjaslice):

@MrNood Bye bye, have a great day!

OpenStudy (mrnood):

btw - that's Grammar Nazi (not Nancy)

OpenStudy (ninjaslice):

@MrNood ;)